# SE4Lip: Speech-Lip Encoder for Talking Head Synthesis to Solve Phoneme-Viseme Alignment Ambiguity

### v1

@@ Abstract
Speech-driven talking head synthesis tasks commonly use general acoustic features (such as HuBERT and DeepSpeech) as guided speech features. However, we discovered that these features suffer from phoneme-viseme alignment ambiguity, which refers to the uncertainty and imprecision in matching phonemes (speech) with visemes (lip). To address this issue, we propose the Speech Encoder for Lip (SE4Lip) to encode lip features from speech directly, aligning speech and lip features in the joint embedding space by a cross-modal alignment framework. The STFT spectrogram with the GRU-based model is designed in SE4Lip to preserve the fine-grained speech features. Experimental results show that SE4Lip achieves state-of-the-art performance in both NeRF and 3DGS rendering models. Its lip sync accuracy improves by 13.7% and 14.2% compared to the best baseline and produces results close to the ground truth videos.
Figure 1: The comparison of the talking head synthesis pipeline when using acoustics features and SE4Lip. The core of SE4Lip is to solve phoneme-viseme alignment ambiguity, which refers to the uncertainty and imprecision in matching phonemes (speech) with visemes (lip). SE4Lip aligns lip features with speech using a cross-modal alignment framework. Thus, SE4Lip can significantly improve the quality of the synthesized videos.
1 Introduction
Talking head synthesis has attracted widespread applications in video conferences [41], film production [18], psychology [8], and other fields. There is an expectation to generate dynamic, realistic, and stable synthetic videos, especially regarding lip movements. The pipeline of talking head synthesis is illustrated in Fig. 1. Speech and visual features are used as conditional inputs to a rendering model, which ultimately synthesizes the video. Since this task is driven by speech, the quality of the synthesized video heavily depends on the quality of speech features.
[…]
This paper proposes SE4Lip to solve the issue of phoneme-viseme alignment ambiguity. SE4Lip is based on the idea of contrastive learning and employs a cross-modal alignment framework to address this issue. SE4Lip models the speech signal using a combination of the STFT spectrogram and the GRU-based model instead of the traditional approach that combines the Mel spectrogram and the CNN-based model [37]. This is because, compared to the STFT spectrogram, the Mel spectrogram compresses the feature space in the frequency domain, leading to the loss of fine-grained frequency information. Additionally, the GRU is more effective in capturing temporal variations, thereby enhancing the ability of speech features to represent lip movements. Inspired by Wav2Lip [30] and SyncNet [7], SE4Lip uses a CNN-based model to extract lip features. The contrastive loss function forces SE4Lip to focus on the causal relationship between phonemes and lip movements rather than linguistic representations. As shown in Fig. 1 with the orange arrows, SE4Lip strengthens the alignment capability of the rendering model, allowing it to synthesize accurate and clear lips. Experimental results show that videos synthesized using SE4Lip outperform those synthesized using acoustics features in different rendering models. We summarize our contributions as follows.
  1. 1.
    Speech-Lip Encoder: To address the phoneme-viseme alignment ambiguity issue, we train a speech encoder specially for talking head synthesis through a cross-modal alignment framework. This framework directly establishes the alignment between speech and lip features in a shared feature space, effectively avoiding the issue of phoneme-viseme alignment ambiguity.
  2. 2.
    Fine-Grained Speech Signal Processing: To preserve fine-grained information in speech features, we propose combining the STFT spectrogram and the GRU model for speech feature extraction. The STFT spectrogram retains more frequency details, while the GRU effectively captures the temporal relationships. This combination of STFT and GRU significantly improves the accuracy and clarity of the lip shapes.
  3. 3.
    Excellent Performance: Compared to four acoustics features, we achieve state-of-the-art results on both NeRF and 3DGS rendering models. Our method achieves a 13.7% improvement in lip sync error confidence and a 14.2% improvement in lip sync error distance compared to the best baseline. The results are also closely aligned with ground truth videos. Furthermore, the results of the ablation experiments validate the effectiveness of the STFT spectrogram and the GRU-based model.
2 Related Work
2.1 Acoustics Feature
Acoustics features aim to extract linguistic representations from speech signals for subsequent tasks such as speech recognition and speaker identification. Representative works include DeepSpeech [1], HuBERT [15], Wav2Vec 2.0 [3] and Whisper [31].
Classification Manner of Pronunciation Examples
@@ 2.1 Acoustics Feature
[…]
Affricate Complete oral closure and slow release of the lung air // & //
@@
    DeepSpeech [1]. DeepSpeech is a speech feature extraction model based on an end-to-end recurrent neural network. This model first combines the CTC loss function with a multi-layer bidirectional LSTM, directly mapping the Mel spectrogram to character sequences.
@@
    HuBERT [15]. HuBERT is a self-supervised learning framework. HuBERT obtains latent acoustic units through downsampling via convolutional layers and uses a BERT-style mask strategy for unit prediction.
@@
    Wav2Vec 2.0 [3]. Wav2Vec 2.0 is a speech feature extraction model based on a convolutional-transformer hybrid architecture. Wav2Vec 2.0 converts raw waveforms into latent vectors using a quantization module combined with contrastive loss for pretraining.
  4. 4.
    Whisper [31]. Whisper employs an encoder-decoder structure to jointly train speech recognition and language identification tasks. Mel spectrograms in Whisper are processed through transformer encoder blocks, which include convolutional layers with GELU [14] activations.
Although these features perform excellently in downstream tasks such as speech recognition and speaker identification, they are not designed with the temporal and frequency detail requirements of the talking head synthesis task in mind. They are designed to extract linguistic representations from the speech signal but cannot enhance the rendering model’s alignment capability.
@@ 2.2 Talking Head Synthesis
Acoustic features are optimized for phoneme classification, focusing on the acoustic distinction between phonemes rather than the alignment relationship between phonemes and visemes. In the pipeline of traditional methods, phoneme-viseme alignment relies solely on the rendering model. However, existing rendering models often align phonemes and visemes through simple feature addition or attention mechanisms, which can not avoid the phoneme-viseme alignment ambiguity.
@@ 3.1 Phoneme-Viseme Alignment Ambiguity
Phoneme-viseme alignment ambiguity refers to the phenomenon that different phonemes can correspond to a similar lip shape. In Tab. 1, we list some phonemes related to phoneme-viseme alignment ambiguity along with their corresponding pronunciation manner. Additionally, we present the specific lip shapes associated with the phoneme-viseme alignment ambiguity phenomenon in Fig. 2. To explain the origin of this phenomenon in detail, we also present a vocal tract anatomical diagram in Fig. 3 from Ramoo [32] to illustrate the organs that affect pronunciation. The larynx or vocal cords are the basis of pronunciation, while the lips and tongue are the articulatory organs. Although the lips are only a part of the speech production system, they are the most intuitive visual feature, which leads to the phoneme-viseme alignment ambiguity issue. For example, as shown in Fig. 4, the lip shapes corresponding to phonemes /d/ and /t/ are similar. However, /d/ is a voiced consonant with strong vocal cord vibrations during pronunciation, while /t/ is a consonant with weak vocal cord vibrations. Despite the significant acoustic differences between /d/ and /t/, the speech features of /d/ and /t/ should be aligned to similar visual features in the cross-modal alignment.
Figure 4: Taking the phonemes /d/ and /t/ as examples to illustrate the phoneme-viseme alignment ambiguity. In the visualization of HuBERT [15] features, there is a noticeable difference between /d/ and /t/. In the visualization of SE4Lip features, there is no significant distinction between /d/ and /t/ because they share similar lip shapes, indicating that SE4Lip differentiates phonemes based on lip shapes rather than acoustic features.
3.2 Drawback of Using Acoustics Features
Acoustics-driven talking head synthesis suffers from the issue of phoneme-viseme alignment ambiguity. The root cause is that acoustics features do not provide additional alignment support in the rendering model, forcing the model to rely solely on its weak alignment, which is achieved through simple feature addition or attention mechanisms. In addition, the goals of acoustics features and talking head synthesis are conflicting. The former aims to extract highly discriminative linguistic representations from speech signals, while the latter seeks to align phonemes with lip shapes. As shown in the scatter plot of Fig. 4, we extracted HuBERT [15] features for /d/ and /t/ and visualized them using PCA [35] and t-SNE [39] methods. It is evident that due to the difference in linguistic representations, HuBERT strongly distinguishes between the two phonemes. However, the talking head synthesis task requires alignment between the phoneme and visual features (particularly the lip shape). Furthermore, acoustics features commonly use the Mel spectrogram as input, leading to the loss of fine-grained frequency information, which is critical for synthesizing lip movements.
Figure 5: Overview of the SE4Lip framework. For a given speech-lip pair, SE4Lip generates embeddings through the corresponding encoder. In the joint embedding space, SE4Lip pulls positive pairs closer and pushes negative pairs farther apart.
@@ 3.2 Drawback of Using Acoustics Features
In general, using acoustics features to drive talking head synthesis results in the following drawbacks: 1) Misalignment of Task Objectives. The purpose of acoustic features is to align speech with linguistic representations, which fundamentally contradicts the requirement of aligning speech with lips in talking head synthesis. This contradiction exacerbates the phoneme-viseme alignment ambiguity, reducing the synthesized video’s lip shape accuracy. 2) Lack of Dynamic Representation. The temporal resolution of acoustics features (typically 25ms frame length) struggles to capture the instantaneous changes in lip movements. For example, the lip closure-opening process for the plosive sound /p/ lasts around 80–120ms. However, the dynamic features of this movement are dropped by the frequency compression of the Mel spectrogram, leading to a blurry lip shape in the synthesis.
@@ 3.3 Contrastive Learning of Speech-Lip
To enhance the alignment capability of the rendering model and address the issue of phoneme-viseme alignment ambiguity, we propose using cross-modal speech features as input for the rendering model. To train such a speech encoder, we use contrastive learning and directly establish the alignment between speech and lip features in a joint embedding space. Our approach forces the speech encoder to focus on the causal relationship between phonemes and visemes rather than speech content or speaker identity. Additionally, we preserve fine-grained speech feature information through detailed processing in both the frequency and time domains. Specifically, we propose using the STFT spectrogram instead of the Mel spectrogram as input to the speech encoder to avoid frequency compression. We also introduce a temporal model to improve the temporal resolution of speech features.
@@ 4.1 Overview
The framework of SE4Lip is shown in Fig. 5. SE4Lip consists of a speech feature extractor and a lip feature extractor. Specifically, in the preprocessing stage, SE4Lip applies STFT transformation to the raw speech to obtain the spectrogram and extract the lip from the raw image. The speech feature extractor employs a GRU-based model [6], while the lip feature extractor uses a CNN-based model. Finally, SE4Lip uses the contrastive loss function to minimize the distance between positive pairs and maximize the distance between negative pairs. With this design, SE4Lip can efficiently capture the dynamic relationship between speech and lip movements. It is worth noting that SE4Lip is trained using continuous speech and image data within a fixed window size, which helps the model capture the dynamic changes in lip movements.
@@
4.2 Speech Feature Extractor
SE4Lip generates the spectrogram using the Short-Time Fourier Transform (STFT). The Mel spectrogram compresses mid-to-high frequency features to simulate the human perception of frequency. Compared to the Mel spectrogram, the frequency distribution of STFT is linear, which preserves more fine-grained frequency information and captures detailed information related to lips. Traditional methods often use the Mel spectrogram, which loses some important frequency information during frequency compression. However, this information is vital for mapping the relationship between speech and lips.

### v2

# Revisiting Speech-Lip Alignment: A Phoneme-Aware Speech Encoder for Robust Talking Head Synthesis

@@ Abstract
Speech-driven talking head synthesis tasks commonly use general acoustic features as guided speech features. However, we discovered that these features suffer from phoneme-viseme alignment ambiguity, which refers to the uncertainty and imprecision in matching phonemes with visemes.To overcome this limitation, we propose a phoneme-aware speech encoder (PASE) that explicitly enforces accurate phoneme–viseme correspondence. PASE first captures fine-grained speech and visual features, then introduces a prediction–reconstruction task to improve robustness under noise and modality absence. Furthermore, a phoneme-level alignment module guided by phoneme embeddings and contrastive learning ensures discriminative audio and visual alignment. Experimental results show that PASE achieves state-of-the-art performance in both NeRF and 3DGS rendering models. Its lip sync accuracy improves by 13.7% and 14.2% compared to the acoustic feature, producing results close to the ground truth videos.
keywords:
Multimodal Alignment , Talking Head , Speech Encoder
journal: Nuclear Physics B
@@
1 Introduction
Figure 1: The comparison of the talking head synthesis pipeline when using acoustic features and PASE. The core of PASE is to solve phoneme-viseme alignment ambiguity, which refers to the uncertainty and imprecision in matching phonemes (speech) with visemes (lip). PASE is an independent encoder that can be seamlessly integrated into various rendering models to enhance the quality of synthesized faces.
Talking head synthesis has attracted widespread applications in video conferences [1], film production [2], psychology [3], and other fields. There is an expectation to generate dynamic, realistic, and stable synthetic videos, especially regarding lip movements. The pipeline of talking head synthesis is illustrated in Figure 1. Speech and visual features are used as conditional inputs to a rendering model, which ultimately synthesizes the video. Since this task is driven by speech, the quality of the synthesized video heavily depends on the quality of the speech features.
Existing work uses general acoustic features as guided speech features. However, we discovered that these features suffer from phoneme-viseme alignment ambiguity. Ambiguity refers to the phenomenon where different phonemes (e.g., /t/ and /d/, /k/ and /g/) correspond to similar visemes (lip). As shown in Figure 1, we first decompose the phoneme–viseme alignment task into two generalizable sub-tasks: 1) extracting visual motion information from acoustic features; 2) mapping the extracted motion information to visual image sequences. As shown by the gray arrows in Figure 1, current work employs acoustic features, such as DeepSpeech [4], HuBERT [5], Wav2Vec 2.0 [6] and Whisper [7], as conditional input to the rendering model for video synthesis. However, these features are designed for tasks like speech recognition or speaker identification and focus on the discriminative power of phonemes in the acoustic domain. Consequently, phoneme-viseme alignment relies on weak alignment within the rendering model, which is simply achieved through methods like feature addition or attention mechanisms. This weak alignment does not adequately address the ambiguity between phonemes and visemes. As a result, several issues can arise when using acoustic features to drive video synthesis: 1) Inaccurate lip shape. Due to the phoneme-viseme alignment ambiguity, the alignment between the lip movements and the phonemes is inaccurate. 2) Blurry lip shape. Acoustic features do not learn the dynamic relationship between the speech signal and lip movements, resulting in a blurry lip shape.
This paper proposes a universal speech encoder, PASE, which can be integrated into various talking head synthesis models to solve the issue of phoneme-viseme alignment ambiguity. PASE introduces a phoneme-aware multimodal alignment framework with contrastive learning. Specifically, PASE models the audio using an STFT spectrogram and a GRU-based encoder, which preserves high-frequency phonetic cues and captures temporal dynamics at the phoneme level. On the visual side, inspired by Wav2Lip [8] and SyncNet [9], PASE employs a multi-channel CNN to extract spatio-temporal lip features. To enhance robustness, a prediction–reconstruction task is applied, where portions of audio or visual features are randomly masked and reconstructed during training. Finally, PASE introduces a phoneme-level alignment module that explicitly enforces the audio–visual correspondence. For each phoneme, audio features serve as anchors and are fused with visual features via a cross-attention mechanism. To further enhance phoneme discrimination, phoneme embeddings are integrated into the audio features before fusion. Contrastive learning is then applied, minimizing the distance between positive audio–visual pairs while maximizing the distance to negative pairs. This design ensures that PASE focuses on phoneme-specific articulatory patterns. As shown in the orange arrows of Figure 1, PASE is designed to function independently of specific rendering models, allowing it to be seamlessly integrated into various backbone architectures to enhance the quality of synthesized faces. Experimental results show that videos synthesized using PASE outperform those synthesized using acoustic features in both NeRF and 3DGS rendering models. We summarize our contributions as follows.
@@ 1 Introduction
    Universal Speech-Lip Encoder: To address the phoneme-viseme alignment ambiguity in talking head synthesis, we propose a universal speech-lip encoder that can be seamlessly integrated into various synthesis models. This encoder directly establishes alignment between speech and lip features within a shared feature space, enabling the extracted speech features to carry rich visual information.
@@
    Phoneme-Level Alignment Framework: We propose a phoneme-level alignment framework that integrates phoneme embeddings into the audio features and employs cross-attention with visual features. By optimizing with contrastive learning at the phoneme granularity, this framework explicitly enforces discriminative correspondence between phonemes and visemes, significantly reducing alignment ambiguity.
@@
    Excellent Performance: Compared to four acoustic features, we achieve state-of-the-art results on both NeRF and 3DGS rendering models. Our method achieves a 13.7% improvement in lip sync error confidence and a 14.2% improvement in lip sync error distance compared to the best baseline. The results are also closely aligned with ground truth videos. Furthermore, the results of the ablation experiments validate the effectiveness of the STFT spectrogram and the GRU-based model.
2 Related Work
2.1 Acoustic Feature
Acoustic features aim to extract linguistic representations from speech signals for subsequent tasks such as speech recognition and speaker identification. Representative works include DeepSpeech [4], HuBERT [5], Wav2Vec 2.0 [6] and Whisper [7]. Although these features perform excellently in downstream tasks such as speech recognition and speaker identification, they are not designed with the temporal and frequency detail requirements of the talking head synthesis task in mind. They are designed to extract linguistic representations from the speech signal but cannot enhance the rendering model's alignment capability.
@@ 2.2 Talking Head Synthesis
Classification Manner of Pronunciation Examples
Plosive / Stop Complete closure in the mouth and sudden release of lung air through the mouth /p/&/b/, /t/&/d/, /k/&/g/
Nasal Complete oral closure in the mouth, the air escapes through the nose /m/&/n/&/N/
Fricative Narrowing with audible friction, close approximation /f/&/v/, /s/&/z/, /T/&/D/, /S/&/Z/
Affricate Complete oral closure and slow release of the lung air // & //
Acoustic features are optimized for phoneme classification, focusing on the acoustic distinction between phonemes rather than the alignment between phonemes and visemes. In the pipeline of traditional methods, phoneme-viseme alignment relies solely on the rendering model. However, existing rendering models often align phonemes and visemes through simple feature addition or attention mechanisms, which can not avoid the phoneme-viseme alignment ambiguity.
@@ 3 Motivation
Figure 4: Taking the phonemes /d/ and /t/ as examples to illustrate the phoneme-viseme alignment ambiguity. In the visualization of HuBERT [5] features, there is a noticeable difference between /d/ and /t/. In the visualization of PASE features, there is no significant distinction between /d/ and /t/ because they share similar lip shapes, indicating that PASE differentiates phonemes based on lip shapes rather than acoustic features.
@@ 3.1 Phoneme-Viseme Alignment Ambiguity
Phoneme-viseme alignment ambiguity refers to the phenomenon that different phonemes can correspond to a similar lip shape. In Table 1, we list some phonemes related to phoneme-viseme alignment ambiguity along with their corresponding pronunciation manner. Additionally, we present the specific lip shapes associated with the phoneme-viseme alignment ambiguity phenomenon in Figure 3. To explain the origin of this phenomenon in detail, we also present a vocal tract anatomical diagram in Figure 3 from Ramoo [38] to illustrate the organs that affect pronunciation. The larynx or vocal cords are the basis of pronunciation, while the lips and tongue are the articulatory organs. Although the lips are only a part of the speech production system, they are the most intuitive visual feature, which leads to the phoneme-viseme alignment ambiguity issue. For example, as shown in Figure 4, the lip shapes corresponding to phonemes /d/ and /t/ are similar. However, /d/ is a voiced consonant with strong vocal cord vibrations during pronunciation, while /t/ is a consonant with weak vocal cord vibrations. Despite the significant acoustic differences between /d/ and /t/, the speech features of /d/ and /t/ should be aligned to similar visual features in the cross-modal alignment.
3.2 Drawback of Using Acoustic Features
@@
Acoustic-driven talking head synthesis suffers from the issue of phoneme-viseme alignment ambiguity. The root cause lies in two drawbacks: 1) Misalignment of Task Objectives. Acoustic features aim to extract highly discriminative linguistic representations, which directly conflict with the synthesis goal of aligning phonemes and lip shapes. As shown in the scatter plot of Figure 4, we extracted HuBERT [5] features for /d/ and /t/ and visualized them using PCA [39] and t-SNE [40] methods. It is evident that due to the difference in linguistic representations, HuBERT strongly distinguishes between the two phonemes. However, the talking head synthesis task requires alignment between the phoneme and visual features (particularly the lip shape). This mismatch forces rendering models to rely on weak alignment mechanisms like feature addition or attention, exacerbating ambiguity and reducing lip-shape accuracy. 2) Lack of Dynamic Representation. The temporal resolution of acoustic features (typically 25ms frame length) struggles to capture the instantaneous changes in lip movements. For example, the lip closure-opening process for the plosive sound /p/ lasts around 80–120ms. However, the dynamic features of this movement are dropped by the Mel spectrogram in the high frequency ranges, leading to a blurry lip shape in the synthesis.
@@ 3.3 Contrastive Learning of Speech-Lip
To enhance the alignment capability of the rendering model and address the issue of phoneme-viseme alignment ambiguity, we propose using cross-modal speech features as input for the rendering model. To train such a speech encoder, we use contrastive learning and directly establish the alignment between speech and lip features in a joint embedding space. Our approach forces the speech encoder to focus on the causal relationship between phonemes and visemes rather than speech content or speaker identity. Additionally, we preserve fine-grained speech feature information through detailed processing in both the frequency and time domains. Specifically, we propose using the STFT spectrogram instead of the Mel spectrogram as input to the speech encoder to avoid the loss of high frequency ranges. We also introduce a temporal model to improve the temporal resolution of speech features.
@@ 4.1 Overview
The framework of PASE is illustrated in Figure 5, which consists of the Audio Encoder, the Visual Encoder, the Prediction and Reconstruction module, and the Phoneme Level Alignment module. During the data preprocessing stage, we first collect the raw images and speech segments corresponding to the same phoneme interval. We extract lip images from the video on a frame-by-frame level, and convert the corresponding speech segments into STFT spectrograms. In the Audio Encoder, we employ a gated recurrent unit (GRU) as the temporal model to extract audio features, thereby capturing critical sequential information from audio. In the Visual Encoder, we design a multi-channel convolutional neural network (CNN) to extract lip-related visual features. The multi-channel CNN is capable of simultaneously capturing both temporal variations and spatial characteristics of lip movements. To enhance robustness, we apply the Prediction and Reconstruction module to both audio and visual features. This module randomly masks a portion of the audio or visual features and requires the model to predict and reconstruct the masked parts using the remaining unmasked features. In the Phoneme Level Alignment module, we treat the audio features as anchors and perform feature fusion through a cross-attention mechanism. During feature fusion, phoneme embeddings are incorporated to enhance the model’s ability to recognize and discriminate specific phonemes. Finally, a contrastive loss is employed to minimize the distance between positive pairs while maximizing the distance between negative pairs, thereby achieving phoneme-level multimodal feature alignment.
Figure 5: The framework of PASE.
4.2 Audio Encoder
Most existing methods follow the traditional paradigm by directly using the Mel spectrogram, without thoroughly investigating their suitability for phoneme-viseme alignment tasks. We argue that the STFT spectrogram is a more appropriate choice. The Mel spectrogram is originally designed to mimic human auditory perception by preserving formant structures (typically within 0.3–4kHz) that convey semantic information, while suppressing high-frequency components to improve efficiency and robustness in speech recognition. However, high-frequency components (>4kHz) carry important articulatory motion cues. For instance, fricatives like /S/ and /s/ exhibit distinct energy distributions in the high-frequency. However, the high-frequency suppression in the Mel spectrogram undermines the model’s ability to capture dynamic visual cues, thereby increasing the alignment ambiguity. Furthermore, the superiority of STFT is also validated through the ablation studies 5.5.
[…]
$z_{t}$ $=\sigma(W_{z}\cdot[h_{t-1},x_{t}])$ (1)
$r_{t}$ $=\sigma(W_{r}\cdot[h_{t-1},x_{t}])$
Here, $\sigma$ denotes the Sigmoid function, $h_{t-1}$ is the hidden state at the previous time step, and $W_{z}$ and $W_{r}$ are the corresponding weight matrices. The computation process of GRU at time step $t$ is represented by Eq. 2.
$\tilde{h}_{t}$ $=\tanh(W_{h}\cdot[r_{t}\odot h_{t-1},x_{t}])$ (2)
$h_{t}$ $=(1-z_{t})\odot h_{t-1}+z_{t}\odot\tilde{h}_{t}$
4.3 Visual Encoder
@@
We aim to provide accurate lip features for aligning speech and lip shapes. Inspired by works [8, 9], we designed a multi-channel CNN-based model to extract lip features. The specific architecture of this model can be found in Table 2. The model captures fine-grained spatial features of the lip images layer by layer through multiple convolutional layers. After processing the lip shape image through the CNN, an embedding is obtained with the same dimension as the speech features, representing the dynamic visual features of the lip shapes.
4.4 Prediction and Reconstruction
To enhance the robustness of the model against feature noise, we introduce a Prediction and Reconstruction task. For the visual features $V\in\mathbb{R}^{T\times D}$ and audio feature sequence $A\in\mathbb{R}^{T\times D}$ belonging to the same phoneme segment, we generate separate binary masking matrices. These matrices randomly select a subset of time steps, where the corresponding features are replaced with learnable vectors. The masking operations for the visual sequence $V$ and the audio sequence $A$ are performed independently and randomly, meaning that one modality, or both simultaneously, can be masked. During training, the model is required to predict and reconstruct the masked segments, thereby providing additional supervisory signals for cross-modal representation learning. The reconstruction loss is defined as Eq. 3.
$\mathcal{L}_{rec}=\lvert\lvert V_{masked}-V_{orig}\rvert\rvert_{2}^{2}+\lvert\lvert A_{masked}-A_{orig}\rvert\rvert_{2}^{2}$ (3)
This auxiliary loss branch effectively alleviates overfitting and enhances the model’s robustness when facing occlusions or modality absence.
4.5 Phoneme Level Alignment
To improve the alignment capability of the model, we propose a phoneme-level alignment strategy. Specifically, for a given phoneme $p$, we use the corresponding audio feature $A_{p}$ as the anchor and apply a cross-attention mechanism to fuse positive and negative visual features. During the fusion process, the phoneme label is embedded to obtain the phoneme label feature $F_{p}$. As formulated in Eq. [q_vector], the phoneme label feature is integrated with the audio feature to generate the query vector $Q_{p}$.
$Q_{p}=A_{p}+F_{p}$ (4)
This design enhances the model’s phoneme-awareness during feature fusion. Subsequently, the visual features $V_{p}$ are used as both key and value vectors, and the fused feature $F_{fusion}$ is obtained according to Eq. 5.
$F_{fusion}=CrossAttn(Q_{p},V_{p},V_{p})$ (5)
For the positive pair $(A_{p},F^{+}{fusion})$ and negative pairs $(A_{p},F^{-}{fusion})$, we employ a contrastive loss, as defined in Eq. 6, to constrain the alignment.
$\mathcal{L}_{\text{con}}=-\log\frac{\exp\left(s(A_{p},F^{+}_{fusion})/\tau\right)}{\exp\left(s(A_{p},F^{+}_{fusion})/\tau\right)+\sum\limits_{F^{-}_{fusion}}\exp\left(s(A_{p},F^{-}_{fusion})/\tau\right)}$ (6)
Here, $s(\cdot)$ denotes the similarity function, and $\tau$ represents the temperature coefficient. Finally, the overall objective function is defined in Eq. 7.
$\mathcal{L}=\mathcal{L}_{\text{con}}+\alpha\cdot\mathcal{L}_{rec}$ (7)
In this formulation, $\alpha$ is a weighting factor that balances the reconstruction loss.
5 Experiments